# EBC-ZIP: Improving Blockwise Crowd Counting with Zero-Inflated Poisson Regression

### v1

@@ IV-B EBC-ZIP-B Compared with State-of-the-Art
TABLE VI: Comparison of our lightweight models with state-of-the-art lightweight approaches on ShanghaiTech A & B, and UCF-QNRF. Models are grouped by parameter size: $<1$ M, $1-10$ M, $10-20$ M, $\geq 20$ M. FLOPs are measured on inputs of size $1920\times 1080$. The best result in each group is highlighted, the overall best result is shown in bold, and the second-best is underlined.
@@
TABLE VII: Inference speed (frames per second, FPS) of the lightweight variants on ShanghaiTech B ($1024\times 768$ resolution). Results are reported on four hardware platforms.
IV-C Comparison with Lightweight Models
To investigate the performance-efficiency trade-off, we compare our proposed lightweight models (EBC-ZIP-P, EBC-ZIP-N, EBC-ZIP-T, and EBC-ZIP-S) with a range of representative compact crowd counting methods. The backbones of these variants can be found in Table I. We divide models into four groups based on parameter size: (1) ultra-lightweight models ($<$1M), (2) lightweight models (1–10M), (3) mid-size models (10–20M), and (4) larger lightweight models ($\geq$20M). Table VI presents the details.
Group I: Ultra-lightweight ($<$1M).
In this group, EBC-ZIP-P achieves the best performance across all three benchmarks, with notably lower MAE on ShanghaiTech B (8.23) and UCF-QNRF (96.29) than LSANet [25], despite having over 50% fewer FLOPs. This demonstrates our model’s superior design in extremely constrained settings.
Group II: Lightweight (1–10M).
EBC-ZIP-N significantly outperforms other methods such as MobileCount [16] and LMSFFNet-XS [26] across all datasets. It achieves the lowest MAE in this group on ShanghaiTech A (60.21), ShanghaiTech B (7.74), and UCF-QNRF (86.46).
Group III: Mid-size (10–20M).
EBC-ZIP-T leads this group with the best results on all benchmarks. It achieves group-best scores of 59.07 MAE on ShanghaiTech A, 6.67 on ShanghaiTech B, and 76.02 on UCF-QNRF, significantly outperforming MobileCount×$\times$2 [16] with fewer parameters and FLOPs.
Group IV: Larger lightweight ($\geq$20M).
EBC-ZIP-S sets a new benchmark in this group and across all lightweight models, achieving 57.20 MAE on ShanghaiTech A, 5.88 on ShanghaiTech B, and 73.32 on UCF-QNRF. These scores represent the best results among all compared lightweight models, and they also approach or surpass those of regular-sized models presented in Table IV.
Across all groups, our EBC-ZIP variants consistently outperform existing methods with comparable or fewer parameters and FLOPs. These results validate the effectiveness and scalability of our EBC-ZIP framework, making it suitable for real-world deployments under various computational constraints.
@@ IV-C Comparison with Lightweight Models
To assess the practical deployment potential of our lightweight models, we measure the average inference speed of each EBC-ZIP variant across four representative hardware platforms: Apple M1 Pro CPU (6 performance + 2 efficiency cores), Apple M1 Pro 14-core GPU, AMD Ryzen 9 5900X CPU (12-core), and NVIDIA RTX 3090 GPU. All experiments are conducted on the ShanghaiTech B dataset using full-precision inference at an input resolution of $1024\times 768$, which reflects real-world surveillance camera scenarios. The PyTorch version is 2.7.1. Table VII reports the average frames per second (FPS) for each model. Our smallest variant, EBC-ZIP-P, achieves the highest throughput on all platforms, reaching 122.22 FPS on AMD 5900X and over 1,000 FPS on RTX 3090, making it ideal for resource-constrained or real-time applications. On the other end, EBC-ZIP-S delivers the best accuracy (MAE = 5.88) but at a higher computational cost, with CPU-side inference limited to under 5 FPS, making it suitable for accuracy-critical but latency-tolerant use cases. EBC-ZIP-N and EBC-ZIP-T offer compelling accuracy-efficiency trade-offs. For example, EBC-ZIP-T achieves 18.37 FPS on AMD 5900X and 249.73 FPS on RTX 3090, while maintaining a competitive MAE of 6.67. These results indicate that mid-sized variants can enable near real-time performance even on high-end CPUs without requiring dedicated GPUs. Overall, our EBC-ZIP family provides a scalable solution space, allowing deployment under varying hardware and latency constraints without compromising on count accuracy.
Backbone: VGG19 Training Resolution
448 672 896
Block Size 16 78.52 75.06 77.92
32 75.73 73.02 75.69
TABLE VIII: Ablation study of training resolution and block size on UCF-QNRF with the VGG19-based backbone. Performances are measured by MAE.
IV-D Influence of Training Resolution and Block Size
We empirically justify our choice of a training resolution of 672 pixels and a block size of 32 pixels on UCF-QNRF through a two-factor ablation study with a VGG-19 backbone. The resulting MAE values are reported in Table VIII. A resolution of 672 pixels consistently outperforms 448 pixels and 896 pixels for both block sizes, and a larger block size of 32 pixels yields lower errors than 16 pixels at every resolution. The combination (672 pixels, 32 pixels) attains the overall best MAE of 73.02, and we therefore adopt this setting for all subsequent UCF-QNRF experiments.
IV-E Visualization
(a) Input Image.
(b) GT Density Map
(c) Structural Zero Map
(d) Predicted Density Map
Figure 6: Qualitative example of EBC-ZIP-B on a ShanghaiTech B test image. The structural-zero map (red = high probability of background/body, blue = head-center region) automatically suppresses background and non-head body areas while preserving only the small central region around each head annotation, resulting in a more accurate density estimate.
Figure 6 presents a moderate-density street-view scene from the test set of ShanghaiTech B together with the corresponding EBC-ZIP outputs. The structural-zero map in 6(c) predicts nearly all non-head pixels (including walls, torso regions, and pavement) as zeros, while preserving only the compact blue blobs centered on annotated heads. This confirms that the $\boldsymbol{\pi}$-head behaves as an implicit, fine-grained foreground segmenter. Leveraging this mask, the predicted density in Fig.6(d) cleanly localizes every head and reproduces the ground-truth count (31) with virtually no background false positives. These results demonstrate that explicit modeling of structural zeros directs the network’s attention to true head centers and improves counting accuracy.
V Conclusion
We presented EBC-ZIP, a crowd-counting framework that improves the Enhanced Blockwise Classification (EBC) framework by integrating zero-inflated Poisson (ZIP) regression. By learning a structural-zero branch to suppress background and non-head body regions, and a Poisson-rate branch to model true head centers, the network tackles the extreme sparsity of density maps without any explicit segmentation supervision. Comprehensive experiments show that the full model, EBC-ZIP-B, establishes new state-of-the-art MAE on ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd, and lowers NAE on the NWPU test set by 21.8% relative to the best previous method. We further introduce a family of lightweight variants (from Pico to Small) that retain these accuracy gains under tight computational budgets and consistently outperform existing models of comparable size on all four benchmarks, highlighting both the effectiveness and scalability of our EBC-ZIP framework. Future work will extend EBC-ZIP to video-based crowd counting and explore additional temporal cues.

### v2

@@ Introduction
Table 5: Comparison of our lightweight models with state-of-the-art lightweight approaches on ShanghaiTech A & B, and UCF-QNRF. Models are grouped by parameter size: $<1$ M, $1-10$ M, $10-20$ M, $\geq 20$ M. FLOPs are measured on $1920\times 1080$ resolution. The best result in each group is highlighted, the overall best result is shown in bold.
To investigate the performance-efficiency trade-off, we compare our proposed lightweight models (ZIP-P, ZIP-N, ZIP-T, and ZIP-S) with a range of representative compact crowd counting methods. We divide models into four groups based on parameter size: (1) ultra-lightweight models ($<$1M), (2) lightweight models (1–10M), (3) mid-size models (10–20M), and (4) larger lightweight models ($\geq$20M). Table 5 presents the details.
Ultra-lightweight ($<$1M). In this group, ZIP-P achieves the best performance across all three benchmarks, with notably lower MAE on ShanghaiTech B (8.23) and UCF-QNRF (96.29) than LSANet, despite having over 50% fewer FLOPs. This demonstrates our model’s superior design in extremely constrained settings.
Lightweight (1–10M). ZIP-N significantly outperforms other methods such as MobileCount and LMSFFNet-XS across all datasets. It achieves the lowest MAE in this group on ShanghaiTech A (58.86), ShanghaiTech B (7.74), and UCF-QNRF (86.46).
Mid-size (10–20M). ZIP-T leads this group with the best results on all benchmarks. It achieves group-best scores of 56.36 MAE on ShanghaiTech A, 6.67 on ShanghaiTech B, and 76.02 on UCF-QNRF, significantly outperforming MobileCount×$\times$2 with fewer parameters and FLOPs.
Larger lightweight ($\geq$20M). ZIP-S sets a new benchmark in this group and across all lightweight models, achieving 55.17 MAE on ShanghaiTech A, 5.83 on ShanghaiTech B, and 73.32 on UCF-QNRF. These scores represent the best results among all compared lightweight models, and they also approach or surpass those of regular-sized models presented in Table 3.
Across all groups, our ZIP variants consistently outperform existing methods with comparable or fewer parameters and FLOPs. These results validate the effectiveness and scalability of our ZIP framework, making it suitable for real-world deployments under various computational constraints.
(a) Input Image.
(b) GT Density Map.
(c) Density Map (DMCount).
(d) Density Map (ZIP).
Figure 6: Qualitative comparison of DMCount and ZIP on an image from NWPU-Crowd (val). Both models utilize VGG19-based encode-decoder structure as the backbone.
Visual Comparison
We also evaluate ZIP and DMCount on the same NWPU-Crowd validation image, using an identical VGG-19 encoder–decoder backbone for fair comparison. As shown in Fig. 6, the two density maps differ markedly. DMCount produces elongated, ribbon-like streaks inside each dense block, which is a typical artifact of its transport-based loss. In contrast, ZIP yields compact, circular kernels that more closely resemble the point annotations, preserving the grid structure of the display. As for accuracy, DMCount under-counts the scene by 114 heads (-6.5%), while ZIP misses only 19 heads (-1.1%), achieving a five-fold reduction in absolute error.
Conclusion
We presented ZIP, a crowd counting framework based on zero-inflated Poisson regression. By learning a structural-zero branch to suppress background and non-head body regions, and a Poisson-rate branch to model blockwise counts, the network tackles the extreme sparsity of density maps without any explicit segmentation supervision. Comprehensive experiments show that the full model, ZIP-B, establishes new state-of-the-art MAE and RMSE on four datasets. We further introduce a family of lightweight variants (from Pico to Small) that retain these accuracy gains under tight computational budgets and consistently outperform existing models of comparable size on all four benchmarks, highlighting both the effectiveness and scalability of ZIP.
Supplementary Material
Proof of Theorem 1
For a single block with parameters $\theta\coloneqq(\pi,\lambda)$ lying in a compact set $\Theta\coloneqq[\pi_{\mathrm{min}},\pi_{\mathrm{max}}]\times[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]\subset(0,1)\times(0,\Lambda]$  [^1] In practice, we generate $\pi$ using sigmoid and $\lambda$ using an average of finite bin centers, and thus we have $10<\pi_{\mathrm{min}}\leq\pi_{\mathrm{max}}<$ 1 and $0<\lambda_{\mathrm{min}.}\leq\lambda_{\mathrm{max}}<\infty$, define
$\mathcal{L}_{\theta}(k)$ $=-\log p_{\theta}(Y=k)$
=$=\begin{cases}-\log[\pi+(1-\pi)e^{-\lambda}],\quad&k=0,\\ -\log(1-\pi)-\lambda+k\log\lambda+\log k!,&k\geq 1.\end{cases}$
Then, for $k\in[0,\lambda_{\mathrm{max}}-1]$, we have the $L$-steep bound:
$|\mathcal{L}_{\theta}(k+1)-\mathcal{L}_{\theta}(k)|$ $=\left|\log\frac{p_{\theta}(k)}{p_{\theta}(k+1)}\right|$
$=\left|\log\frac{k+1}{\lambda}\right|$
$\leq\sup_{k,\lambda\in[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]}\left|\log\frac{k+1}{\lambda}\right|$
$=\log\frac{k_{\mathrm{max}}+1}{\lambda_{\mathrm{min}}}\eqqcolon L,$
where $k_{\mathrm{max}}$ is the largest count that ever appears in one block in the training set. Note that $k_{\mathrm{max}}$ is restricted to a relatively small number to suppress label noise following EBC framework (Ma, Sanchez, and Guha 2024).
Following (Wang et al. 2020a), we flatten the spatial dimensions to avoid cluttered notations. For a full image with $n\coloneqq hw$ blocks, let
$g(\boldsymbol{z};\boldsymbol{\theta})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{\boldsymbol{\theta}_{i}}(\boldsymbol{z}_{i}),$
For two annotation-prediction pairs $(\boldsymbol{z},\boldsymbol{\theta})$ and $(\boldsymbol{z}^{\prime},\boldsymbol{\theta})^{\prime}$:
$|g(\boldsymbol{z};\boldsymbol{\theta})-g(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})|$
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}})-\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})\right|+\frac{1}{n}\sum_{i=1}^{n}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})-\mathcal{L}_{\boldsymbol{\theta}_{i}^{\prime}}({\boldsymbol{z}_{i}^{\prime}})\right|$
$\leq$ $\frac{L}{n}\sum_{i=1}^{n}\left|\boldsymbol{z}_{i}-\boldsymbol{z}_{i}^{\prime}\right|+M\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1},$
where
$M\coloneqq\sup_{k,\theta}\|\nabla_{\theta}\mathcal{L}_{\theta}(k)\|<\infty$
due to the compactness of $\Theta$.
Combining the two sums gives
$|g(\boldsymbol{z};\boldsymbol{\theta})-g(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})|$ $\leq L\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\|_{1}+M\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1}$
$\leq C\|(\boldsymbol{z},\boldsymbol{\theta})-(\boldsymbol{z}^{\prime},\boldsymbol{\theta}^{\prime})\|_{1}$
where $C\coloneqq\max(L,M)<\infty$. Hence, the ZIP loss is $C$-Lipschitz under the joint $\mathcal{L}_{1}$ norm.
As in the supplement of (Wang et al. 2020a), we treat the crowd counting model as $n$ identical scalar regressors. Lemma E (b) in (Wang et al. 2020a) gives
$\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathrm{ZIP}}\circ\mathcal{F}^{\prime})\leq C\cdot n\cdot\mathcal{R}_{\mathcal{S}}(\mathcal{H}).$
Plugging this into Theorem 26.5 in (Shalev-Shwartz and Ben-David 2014) yields, $\forall\delta\in(0,1)$, with the probability at least $1-\delta$ over the draw of the training sample $\mathcal{S}$ of size $K$,
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{ZIP}},\mathcal{L}_{\mathrm{ZIP}})\leq$ $\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{ZIP}})+2\cdot C\cdot n\cdot R_{\mathcal{S}}(\mathcal{H})+$
$5B\sqrt{\frac{2\log(8/\delta)}{K}},$
where $B=\sup_{\boldsymbol{X},\boldsymbol{z}}\mathcal{L}_{\mathrm{ZIP}}(\boldsymbol{z},f(\boldsymbol{X}))$. ∎
Inference Speed
@@ Conclusion
Table 6: Inference speed (frames per second, FPS) of the lightweight variants on ShanghaiTech B ($1024\times 768$ resolution). Results are reported on four hardware platforms.
@@
To assess the practical deployment potential of our lightweight models, we measure the average inference speed of each ZIP variant across four representative hardware platforms: Apple M1 Pro CPU (6 performance + 2 efficiency cores), Apple M1 Pro 14-core GPU, AMD Ryzen 9 5900X CPU (12-core), and NVIDIA RTX 3090 GPU. All experiments are conducted on the ShanghaiTech B dataset using full-precision inference at an input resolution of $1024\times 768$, which reflects real-world surveillance camera scenarios. The PyTorch version is 2.7.1. Table 6 reports the average frames per second (FPS) for each model. Our smallest variant, ZIP-P, achieves the highest throughput on all platforms, reaching 122.22 FPS on AMD 5900X and over 1,000 FPS on RTX 3090, making it ideal for resource-constrained or real-time applications. On the other end, ZIP-S delivers the best accuracy (MAE = 5.83) but at a higher computational cost, with CPU-side inference limited to under 5 FPS, making it suitable for accuracy-critical but latency-tolerant use cases. ZIP-N and ZIP-T offer compelling accuracy-efficiency trade-offs. For example, ZIP-T achieves 18.37 FPS on AMD 5900X and 249.73 FPS on RTX 3090, while maintaining a competitive MAE of 6.67. These results indicate that mid-sized variants can enable near real-time performance even on high-end CPUs without requiring dedicated GPUs. Overall, our ZIP family provides a scalable solution space, allowing deployment under varying hardware and latency constraints without compromising on count accuracy.
Method ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU-Crowd (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
BL (Ma et al. 2019) 62.8 101.8 7.7 12.7 88.7 154.8 93.6 470.3
DMCount (Wang et al. 2020a) 59.7 95.7 7.4 11.8 85.6 148.3 70.5 357.6
EBC (Ma, Sanchez, and Guha 2024) 62.3 98.9 7.0 10.9 77.2 130.4 39.6 95.8
ZIP (ours) 58.7 93.6 6.3 10.3 73.0 126.1 38.3 90.3
Improvement over EBC 5.9% 2.9% 10.0% 5.5% 5.4% 3.2% 3.2% 5.7%
Table 7: Comparison of ZIP with other frameworks on ShanghaiTech A & B, UCF-QNRF and NWPU-Crowd. All frameworks use the VGG19-based backbone. Results demonstrate that our framework ZIP significantly outperforms other methods, achieving the lowest MAE and RMSE on all four benchmarks. Our ZIP also consistently improves the EBC framework, at most by 10.0%. The best results are highlighted in bold font, and the second best results are underscored.
Suffix Size (M) ShanghaiTech A ShanghaiTech B UCF-QNRF
MAE RMSE NAE (%) MAE RMSE NAE (%) MAE RMSE NAE (%)
-P 0.81 71.18 109.60 16.69 8.23 12.62 6.98 96.29 161.82 14.40
-N 3.36 58.86 94.63 14.15 7.74 12.14 6.33 86.46 147.64 12.60
-T 10.53 56.36 90.55 13.26 6.67 9.90 5.52 76.02 129.40 11.10
-S 33.60 55.17 88.99 11.97 5.83 9.21 4.58 73.32 125.09 10.40
-B 105.60 47.81 75.04 11.06 5.51 8.63 4.48 69.46 121.88 10.18
Table 8: Comparison of all ZIP variants on ShanghaiTech A & B, and UCF-QNRF. Best results are shown in the bold typeface and the second best are underlined.
Framework Comparison
We compare ZIP framework with several existing methods on the four benchmarks, shown in Table 7. Our approach sets a new state-of-the-art on ShanghaiTech B, UCF-QNRF and NWPU-Crowd. It also achieves the lowest MAE and the second lowest RMSE on ShanghaiTech A. Notably, when compared directly with its EBC baseline, our ZIP model demonstrates significant performance gains across all datasets, reducing the MAE by a margin of 3.2% to 10.0%. These improvements validate the efficacy of incorporating zero-inflated Poisson regression for crowd counting.
Comparison of All ZIP Variants
Table 8 reports the full results of the five ZIP variants: Pico (-P), Nano (-N), Tiny (-T), Small (-S) and Base (-B), on ShanghaiTech A, ShanghaiTech B and UCF-QNRF. The variants differ only in backbone capacity (from 0.8 M to 105 M parameters); all heads, losses and training settings are identical. From this table, we can observe three trends:
  • Monotonic accuracy gain with scale. From -P to -B every step produces lower error on all datasets. MAE on ShanghaiTech A drops from 71.2 to 47.8 (-32.8%), while UCF-QNRF MAE falls from 96.3 to 69.5 (-27.8%). A similar monotone decrease is observed for RMSE and NAE.
  • Diminishing returns. The relative improvement per additional million parameters shrinks after the -T model (10 M). For example, moving from -T to -S cuts ShanghaiTech A MAE by 3.7% whereas the jump from -N to -T yields 1.7$\times$ that reduction with fewer added parameters. This suggests that -T or -S may offer the best accuracy–efficiency trade-off for many deployments.
  • Consistent generalization across datasets. Ranking of variants is identical on ShanghaiTech A/B and UCF-QNRF, indicating that gains are not dataset-specific but stem from the scalable ZIP formulation itself.
Overall, the results demonstrate that ZIP scales gracefully: the smallest model remains competitive for edge devices, whereas the largest model achieves state-of-the-art accuracy.
Figure 7: Examples of annotation noise in NWPU-Crowd. Left column: input images with zoom-in views of the regions outlined in red. Center column: ground-truth density map with total count. Right column: density map with total count predicted by ZIP-B. Top row: mirror reflections labeled as real people. Bottom row: photo-realistic audience on a display screen left unlabeled.
Annotation Noise in NWPU-Crowd
Although NWPU-Crowd is currently the largest high-resolution benchmark, we observed non-negligible label noise. Fig. 7 presents two typical failure modes: over-labeled reflections (top row) and under-labeled screen display (bottom row).
In the top row, the red box highlights strong reflections of pedestrians on a glass facade. These mirror images have been annotated as real heads, yet they are visually less consistent with surrounding people, resulting a total count of 439. ZIP-B correctly ignores the reflections, producing a lower but more realistic count of 272.
The bottom row depicts a conference hall is shown on a large display. The audience visible inside the screen looks photo-realistic, but none of these heads are annotated (GT = 221). The model nevertheless detects them, yielding a higher prediction (671). Qualitatively, the additional detections correspond to the people shown on the screen.
Figure 8: Two typical failure cases of ZIP-B. Left: input image with a zoom-in of the critical region (red box). Center: predicted density map with total count. Right: ground-truth density map with total count.
These examples underline two points: (i) reported errors can stem from annotation artifacts rather than model failure, and (ii) the structural-zero + ZIP formulation is robust enough to down-weight unrealistic cues (mirror images) while still detecting plausible but unlabeled heads (video feed). Future benchmarks would benefit from more consistent annotation protocols or confidence tags to flag uncertain regions.
Failure Analysis
Despite its overall strong performance, ZIP-B still fails under certain visual ambiguities. Fig. 8 highlights two typical errors. The top row corresponds to false positive on man-shaped artifact: in an indoor museum scene, a post on the rope barrier strongly resembles a standing person in both shape and color, and as a result, ZIP-B misclassifies the post as a human. In the bottom row, rows of soldiers wear camouflage that blends with the background trucks. Although nearby soldiers are detected reliably, the model fails to pick up many small, low-contrast heads in the far field, leading to an under-count of -94. These failure cases suggest that incorporating additional sensing modalities, particularly infrared imagery, may help suppress person-like artifacts and recover low-contrast targets, thereby reducing both false positives and false negatives. Exploring how the ZIP formulation can be extended to multi-modal crowd counting will be an important direction for our future work.
(a) Block size = 8.
(b) Block size = 16.
(c) Block size = 32.
Figure 9: Blockwise count distributions across common crowd counting benchmarks at different block sizes (8, 16, 32). For each dataset, we first generate ground-truth pixel-level density maps from point annotations, without using Gaussian smoothing. Then, we slide a $B{\times}B$ window with stride 1 across the map, summing the density within the window to obtain its count. Bars show the percentage of windows with count values 0,1,2,3,40,1,2,3$\geq$ 40 , 1 , 2 , 3 , ≥ 4 (log-scale $y$-axis).
Blockwise Count Distribution
For each dataset’s training split, we form a pixel-level density map by placing 1 at each annotated head location and 0 elsewhere, without Gaussian smoothing being applied. We then slide an axis-aligned $B{\times}B$ window over the map with stride 1 pixel and record the integer count equal to the sum of densities in that window (equivalently, the number of annotated heads whose pixel coordinates fall inside the window). We repeat for $B\in\{8,16,32\}$ and aggregate counts over all windows from all images, converting to percentages. Fig. 9 illustrates the results. As the block size (window size) $B$ increases, zero percentages drop because each window integrates over a larger receptive field and is therefore more likely to include at least one head annotation. However, even at $B{=}32$ the distributions remain heavily zero-inflated. For instance, $>80\%$ zeros in UCF-QNRF and $>90\%$ in ShanghaiTech B and NWPU-Crowd. This indicates that annotation sparsity is a persistent regime rather than an artifact of using very fine windows. Note that increasing $B$ also coarsens supervision: spatial detail about where heads occur within a window is lost, making it harder for a crowd counting model to localize density peaks and learn fine-grained cues. In the extreme limit, $B$ equal to the entire image collapses supervision to a single global count label; the learning problem degenerates to (image-level) weakly supervised counting with no spatial guidance at all. Our ZIP formulation is motivated by operating in the practically useful regime of relatively small block size, where supervision is highly sparse yet still spatially informative, necessitating an explicit treatment of excess zeros.
Additional Qualitative Visualizations
We curate ten images from the validation split of NWPU-Crowd to demonstrate the qualitative behavior of ZIP-B in five representative scenes (two images per scene). For clarity we organize the examples into a regular group (background-only, sparse, and crowded) and a challenging group (multi-scale and low-illumination). The results are shown in Fig. 10 and Fig. 11, respectively.
Background-only (Fig. 10(a)). The structural-zero branch assigns a uniformly high zero-probability to every pixel, and the density branch outputs all-zero maps, matching the ground-truth counts of 0.
Sparse (Fig. 10(b)). With only a handful of pedestrians, the structural-zero map suppresses the background while isolating small blue blobs around true head centers; the resulting density maps reproduce the exact person counts.
Crowded (Fig. 10(c)). In the two stadium-like scenes containing several thousand people, the model still produces a clean, high-resolution density map. The total count deviates from ground truth by less than 0.5%, confirming that ZIP modeling scales to extreme densities.
Multi-scale (Fig. 11(a)). The foreground heads are large and the background heads extremely small. The structural-zero map adapts its granularity: it keeps broad foreground regions non-zero while aggressively filtering distant background pixels, enabling accurate detection across scales.
Low-illumination (Fig. 11(b)). Despite poor lighting, the structural-zero branch reliably removes dark background regions, allowing the density branch to localize heads and deliver a near-exact total count.
Together, these qualitative results show that the explicit modeling of structural zeros enables ZIP-B to generalize from empty backgrounds to dim, highly congested, and multi-scale scenes without any segmentation supervision.
(a) Background-only instances.
(b) Sparse scenarios.
(c) Highly crowded scenarios.
Figure 10: Visualized results on background-only, sparse and highly congested scenes in NWPU-Crowd (predictions by ZIP-B). Each row shows, from left to right, the input image, the predicted structural-zero map, the predicted density map, and the ground-truth density map. In the background-only examples (Fig. 10(a)), the structural-zero branch correctly classifies the entire image as background, yielding an all-zero density map. In the sparse cases (Fig. 10(b)), ZIP-B still localizes every head and reproduces the exact crowd count. As for the highly crowded scenes (Fig. 10(c)), ZIP-B can still produce accurate density maps and counts.
(a) Multi-scale scenes.
(b) Low-illumination examples.
Figure 11: Qualitative results on two challenging scenarios in NWPU-Crowd (predictions by ZIP-B): multi-scale (Fig. 11(a)) and low-illumination (Fig. 11(b)). For each row, the columns show (1) input image, (2) predicted structural-zero map, (3) predicted density map, and (4) ground-truth density map. In the multi-scale examples (Fig. 11(a)), the scenes contain a large variety of head sizes, ranging from large, close-up people to tiny, distant heads; ZIP-B suppresses background pixels while simultaneously detecting heads at all scales, yielding an accurate total count. In the low-illumination examples (Fig. 11(b)), the structural-zero branch can still filter the background, enabling the density branch to produce a clean, well-localized prediction despite poor lighting.